6
-3 2/5=


7
-1 3/5=

Answer by fractalier(6550)

6 - 3 2/5 = 5 5/5 - 3 2/5 = 2 3/5
and
7 - 1 3/5 = 6 5/5 - 1 3/5 = 5 2/5